if 3 + 9 = 31, 15 + 12 = 45 ,18 + 9 = 36, then 12 +27 is equal to 
  1.  94
  2. 14
  3. 49
  4. 53
  5. 100

Grade:10

4 Answers

Yugabrat Gogoi
9 years ago
ASHWIN
 
ACCORDING TO ME ITS 94.
 
BECAUSE-
IF U NOTICE IT- ALL THE NUMBERS ARE MULTIPLE OF 3
 
IT IS DERIVED THIS WAY
EACH OF THE VARIABLES IN THE NUMBERS ARE DIVISIBLE BY 3. SO, AFTER THEY ARE DIVIDED BY 3
LIKE 3/3 +9/3
15/3 + 12/3
18/3 + 9/3
12/3 +27/3
AND WE GET
1+3
5+4
6+3
4+9
NOW OBSERVE CAREFULLY, REMOVE THE PLUS SIGN AND REVERSE THE DIGITS.
WE GET -
31
45
36
94
THEREFORE,
3+9=31
15+12=45
18+9=36 AND
12+27=94
 
THANKYOU,
 
YUGABRAT GOGOI
Yugabrat Gogoi
9 years ago
PLEASE REPLY ABT U LIKE IT OR NOT.
Ashwin
9 years ago
thx man btw did did you thought upto this extent
Yugabrat Gogoi
9 years ago
yes man